# Geometric structures on the quaternionic unit ball and slice regular Möbius transformations

### v1

@@ 6. Properties of the slice geometric structures on $\mathbb{B}$
For every $I\in\mathbb{S}$, the slice Riemannian metric $G$ of $\mathbb{B}$ restricted to the tangent bundle $T\mathbb{C}_{I}$ of $\mathbb{C}_{I}$ is $g_{I}$. More precisely, we have $G_{q}(\alpha,\beta)=(g_{I})_{q}(\alpha,\beta)$, whenever $q\in\mathbb{D}_{I}$ and $\alpha,\beta\in\mathbb{C}_{I}$. In other words, the Riemannian metric on $\mathbb{D}_{I}$ induced by the slice Riemannian metric $G$ of $\mathbb{B}$ is the usual hyperbolic metric under the canonical identification $\mathbb{D}_{I}\simeq\mathbb{D}$.
@@
It also follows from the previous results that the slice Riemannian metric of $\mathbb{B}$ can be determined by its value at a single slice. This will make use of the non-regular Möbius transformation $C_{u}$ defined by $C_{u}(q)=u^{-1}qu$, where $u\in\mathrm{Sp}(1)$.
@@
Since $C_{u}$ is linear for every $u\in\mathrm{Sp}(1)$, it coincides with its differential. Then, the first formula for $G$ in Theorem 6.1 implies the claimed invariance. From this, the representation formula is just a restatement of the $C_{u^{-1}}$-invariance. ∎
@@
$\Omega_{q}(\alpha,\beta)=\frac{\operatorname{Im}\big{(}(1-q^{2})^{-1}(\alpha-q\alpha q)(\overline{q-q\beta q})(1-\overline{q}^{2})^{-1}\big{)}}{(1-|q|^{2})^{2}},$
for every $q\in\mathbb{B}$ and $\alpha,\beta\in\mathbb{H}$. In particular, it satisfies the following properties.
@@
The expression for $\Omega_{q}$ in the statement is a direct consequence of the definition of $\Omega$ and Theorem 5.7. The ant-symmetry of $\Omega$ follows immediately from the property $\operatorname{Im}(\alpha\overline{\beta})=-\operatorname{Im}(\beta\overline{\alpha})$ for $\alpha,\beta\in\mathbb{H}$. On the other hand, the non-degeneracy of $\Omega$ follows from Definition 5.6, the fact that $\operatorname{d\!}\>(\mathcal{F}_{q})_{q}$ is an isomorphism (see Proposition 3.3) and the non-degeneracy of the bilinear form $(\alpha,\beta)\mapsto\operatorname{Im}(\alpha\overline{\beta})$. Finally, for a given $I\in\mathbb{S}$, the restriction of $\Omega$ to the tangent bundle $T\mathbb{D}_{I}$ is obtained by considering the values
@@
The equivariance formulas from the statement are easily obtained by substitution on the expressions for $H$ and $\Omega$ given in Theorems 5.7 and 6.4. ∎
@@
With the previous results at hand, we compare the $\mathrm{Sp}(1,1)$-invariant metric $\widehat{G}$ and a couple of well known geometric quaternionic structures with our slice geometric ones. For simplicity, we will discuss the $4$-dimensional case which is clearly enough for our purposes. In the first place, we have the notion of a hyper-Kähler structure. On a $4$-dimensional manifold, such structure is given by a Riemannian metric and three integrable complex structures orthogonal for the metric and satisfying the usual quaternionic commutation relations. Further details on their definition and properties can be found in [1]. As stated in Theorem 14.13 of the latter reference, Berger proved that every hyper-Kähkler manifold is Ricci flat. However, the metric $\widehat{G}$ is known to have strictly negative sectional curvature (see [12]) and so strictly negative Ricci curvature. Hence, from Theorem 5.1 it follows that there is no hyper-Kähler structure on $\mathbb{B}$ that is invariant under the $\mathrm{Sp}(1,1)$-action by non-regular Möbius transformations. In second place, there is the notion of quaternion-Kähler manifold. We refer again to [1] for further details. One possible definition, in the $4$-dimensional case, is obtained by requiring the given Riemannian manifold to have holonomy $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$. This is in fact the case for the metric $\widehat{G}$, since the holonomy and the isotropy coincide for Riemannian symmetric spaces (see [12]). Hence, $\mathbb{B}$ with the metric $\widehat{G}$ is quaternion-Kähler. It is also known, as discussed in Section D from Chapter 14 of [1] and in page 92 of [11], that every quaternion-Kähler manifold carries three locally defined almost-complex structures and three locally defined $2$-forms. The latter allow to construct a globally defined $4$-form. However, in the case of $\mathbb{B}$ with the metric $\widehat{G}$ the locally defined objects cannot be extended globally. In particular, the quaternion-Kähler structure obtained from the $\mathrm{Sp}(1,1)$-action does not yield associated globally defined $2$-forms. In conclusion, the $\mathrm{Sp}(1,1)$-action by non-regular Möbius transformations fails to provide a notion of Kähler structure on $\mathbb{B}$ as some sort of $2$-form involving $\mathbb{H}$ that could also be nicely related to a corresponding Riemannian metric. However, the slice regular Kähler structure $\Omega$ does provide a natural $2$-form with the Kähler-like properties obtained in Theorem 6.4. Furthermore, by construction, $\Omega$ exhibits as much invariance under $\mathcal{M}(\mathbb{B})$ as the latter allows considering that it is not a group. Since, $\Omega$ is $\operatorname{Im}(\mathbb{H})$-valued we can always consider the components with respect to an orthonormal basis of $\operatorname{Im}(\mathbb{H})$ to obtain three real-valued $2$-forms. The slice Hermitian metric $H$ and the slice Riemannian metric $G$ have similar properties as stated in Theorems 5.7 and 6.1, respectively, and their corollaries. More precisely, they both exhibit as much invariance as $\mathcal{M}(\mathbb{B})$ allows. Also, a certain degree of invariance for all three slice geometric structures has been obtained in Corollaries 6.3 and 6.7. The latter have provided the sort of Representation Formulas found in the theory of slice regular functions, but now for geometric structures on $\mathbb{B}$ built from slice regular functions. Finally, the slice Riemannian metric $G$ and the slice Kähler form $\Omega$ can be considered nicely related in a similar way as it occurs in the classical complex case. The reason is that $H$ and $\Omega$ are the (quaternionic) real and imaginary parts, respectively, of the slice Hermitian metric $H$. With all three of them built out of the same principle: regular Möbius transformations.
@@ Acknowledgment
This research was partially supported by SNI-Conacyt and by Conacyt Grants 280732 and 61517.

### v2

@@ 6. Properties of the slice geometric structures on $\mathbb{B}$
For every $I\in\mathbb{S}$, the slice Riemannian metric $G$ of $\mathbb{B}$ restricted to the tangent bundle $T\mathbb{C}_{I}$ of $\mathbb{C}_{I}$ is $g_{I}$. More precisely, we have
$G_{q}(\alpha,\beta)=(g_{I})_{q}(\alpha,\beta),$
whenever $q\in\mathbb{D}_{I}$ and $\alpha,\beta\in\mathbb{C}_{I}$. In other words, the Riemannian metric on $\mathbb{D}_{I}$ induced by the slice Riemannian metric $G$ of $\mathbb{B}$ is the usual hyperbolic metric under the canonical identification $\mathbb{D}_{I}\simeq\mathbb{D}$.
As noted in our Introduction, a Riemannian metric on $\mathbb{B}$ has been constructed in [1] following a function space theoretic approach. This was achieved by considering the quaternionic Hardy space on $\mathbb{B}$. We will now restate some of the results from [1, Theorem 1.1] within the setup of our notation. This will allow us to consider the Riemannian metric defined in [1] and compare it with our Riemannian metric $G$. For this, we will discuss some of the notions studied in [1]. We refer to the latter for further details and claims below on this matter.
Let us denote by $H^{2}(\mathbb{B})$, the quaternionic Hardy space, which consists of functions defined by (convergent) power series of the form
$\sum_{n=0}^{\infty}q^{n}a_{n},$
where $(a_{n})_{n}$ is an absolutely square summable quaternionic sequence. The space $H^{2}(\mathbb{B})$ admits the positive definite quaternionic inner product given by
$\bigg{\langle}\sum_{n=0}^{\infty}q^{n}a_{n},\sum_{n=0}^{\infty}q^{n}b_{n},\bigg{\rangle}_{H^{2}(\mathbb{B})}=\sum_{n=0}^{\infty}\overline{b}_{n}a_{n},$
for which $H^{2}(\mathbb{B})$ turns out to be a quaternionic reproducing kernel Hilbert space. Its reproducing kernel is obtained from the following elements of $H^{2}(\mathbb{B})$
$k_{q}(w)=\sum_{n=0}^{\infty}w^{n}\overline{q}^{n},$
where both $q,w\in\mathbb{B}$. Proceeding in a way similar to the complex case, [1] defines a pseudo-hyperbolic distance given by
(6.2) $\delta(p,q)=\Bigg{(}1-\bigg{|}\bigg{\langle}\frac{k_{p}}{\|k_{p}\|_{H^{2}(\mathbb{B})}},\frac{k_{q}}{\|k_{q}\|_{H^{2}(\mathbb{B})}}\bigg{\rangle}_{H^{2}(\mathbb{B})}\bigg{|}^{2}\Bigg{)}^{\frac{1}{2}},$
for every $p,q\in\mathbb{B}$. As the following result states, it is proved in [1] that this distance comes from a Riemannian metric. The next result is in fact a subset of [1, Theorem 1.1] in the setup of our current notation.
Theorem 6.3 (Arcozzi-Sarfatti [1]).
The distance $\delta$ defined in (6.2) is induced by a Riemannian metric $\widetilde{G}$. Furthermore, for every $q\in\mathbb{B}$ and $\alpha\in T_{q}\mathbb{B}$, the norm of $\alpha$ with respect to $\widetilde{G}_{q}$ is given by
$|\alpha|_{q}^{2}=\widetilde{G}_{q}(\alpha,\alpha)=\frac{|\pi_{I}(\alpha)|^{2}}{(1-|q|^{2})^{2}}+\frac{|\pi_{I}^{\perp}(\alpha)|^{2}}{|1-q^{2}|^{2}}.$
The next result proves that our metric $G$ and Arcozzi-Sarfatti’s metric, denoted here by $\widetilde{G}$, turn out to be exactly the same.
Theorem 6.4.
The slice Riemannian metric $G$ of $\mathbb{B}$ satisfies
$\|\alpha\|^{2}_{q}=G_{q}(\alpha,\alpha)=\frac{|\pi_{I}(\alpha)|^{2}}{(1-|q|^{2})^{2}}+\frac{|\pi_{I}^{\perp}(\alpha)|^{2}}{|1-q^{2}|^{2}},$
for every $q\in\mathbb{B}$ and $\alpha\in T_{q}\mathbb{B}$. In particular, $G=\widetilde{G}$, where the latter is the Riemannian metric considered in Theorem 6.3.
Proof.
Once the identity in the statement has been proved, a polarization argument implies that $G=\widetilde{G}$. Hence, we proceed to establish the aforementioned identity.
Let us fix $q\in\mathbb{B}$ and $\alpha\in T_{q}\mathbb{B}$. Note that each tangent space of $\mathbb{B}$ is canonically identified with $\mathbb{H}$, and so we consider $\alpha\in\mathbb{H}$. Let us assume that $q\in\mathbb{C}_{I}$, where $I\in\mathbb{S}$, and let us choose $J\in\mathbb{S}$ which anti-commutes with $I$. Thus, we can consider the decomposition
$\alpha=\alpha_{1}+\alpha_{2}J,$
where $\alpha_{1},\alpha_{2}\in\mathbb{C}_{I}$. Hence, $|\alpha_{1}|^{2}=|\pi_{I}(\alpha)|^{2}$ and we also have
$|\alpha_{2}|^{2}=-\pi_{I}^{\perp}(\alpha)^{2},\quad|\alpha_{2}|^{2}=|\pi_{I}^{\perp}(\alpha)|^{2},$
where the first identity follows from the discussion at the end of the proof of Theorem 6.1, and the second is a consequence of the first. Hence, by using these identities in the last formula for $G_{q}$ in Theorem 6.1 we obtain
$\|\alpha\|^{2}_{q}=G_{q}(\alpha,\alpha)$ $=\frac{|\alpha_{1}|^{2}+|\alpha_{2}|^{2}}{(1-|q|^{2})^{2}}-4\frac{|\operatorname{Im}(q)|^{2}|\alpha_{2}|^{2}}{|1-q^{2}|^{2}(1-|q|^{2})^{2}}$
$=\frac{|\pi_{I}(\alpha)|^{2}+|\pi_{I}^{\perp}(\alpha)|^{2}}{(1-|q|^{2})^{2}}-4\frac{|\operatorname{Im}(q)|^{2}|\pi_{I}^{\perp}(\alpha)|^{2}}{|1-q^{2}|^{2}(1-|q|^{2})^{2}}$
$=\frac{|\pi_{I}(\alpha)|^{2}}{(1-|q|^{2})^{2}}+\bigg{(}\frac{|1-q^{2}|^{2}-4|\operatorname{Im}(q)|^{2}}{(1-|q|^{2})^{2}}\bigg{)}\frac{|\pi_{I}^{\perp}(\alpha)|^{2}}{|1-q^{2}|^{2}}.$
Hence, it is enough to prove that the quantity in parentheses is identically $1$ for every $q\in\mathbb{B}$. This is a simple computation that we show for the sake of completeness. Recall that $q\in\mathbb{C}_{I}$, so we write $q=x+yI$ with $x,y\in\mathbb{R}$. Then, we have
$|1-q^{2}|^{2}-4|\operatorname{Im}(q)|^{2}$ $=|1-(x^{2}-y^{2}+2xyI)|^{2}-4y^{2}$
$=(1-x^{2}+y^{2})^{2}+4x^{2}y^{2}-4y^{2}$
$=1+x^{4}+y^{4}-2x^{2}+2y^{2}-2x^{2}y^{2}+4x^{2}y^{2}-4y^{2}$
$=1+x^{4}+y^{4}-2x^{2}-2y^{2}+2x^{2}y^{2}$
$=(1-x^{2}-y^{2})^{2}$
$=(1-|q|^{2})^{2},$
which thus complete the proof. ∎
Remark 6.5.
A number of properties for the slice Riemannian metric $G$ can be derived from Theorem 6.4 thanks to the developments carried out in [1]. For the slice Riemannian metric $G=\widetilde{G}$, the latter computes the isometry group, the Lipschitz functions, geodesics and induced volume form on the boundary of $\mathbb{B}$, as well as their relationship with regular functions.
On the other hand, it is important to emphasize the difference in techniques used to obtain $G=\widetilde{G}$ in [1] as opposed to our approach. The construction of this metric in [1] follows a function theoretic approach that mirrors the pseudo-hyperbolic distance in the complex unit disk as expressed in terms of a reproducing kernel. Our approach mirrors the complex case as well, which was described in section 4, but we do so from the viewpoint of transformations as geometric objects. Even though the quaternionic case of $\mathbb{B}$ does not entitle a group from regular Möbius transformations, we have been able to extract enough group-like properties to construct the metric $G$ by looking for an invariant geometric tensor. One advantage of our method is the simplicity, in terms of computations needed, to obtain the slice Riemannian metric.
At any rate, Theorem 6.4 ensures that both the function space theory and geometric/group-like approaches to construct a Riemannian metric lead to the same solution: the slice Riemannian metric $G$.
@@
We now prove that the slice Riemannian metric of $\mathbb{B}$ can be determined by its value at a single slice. This will make use of the non-regular Möbius transformation $C_{u}$ defined by $C_{u}(q)=u^{-1}qu$, where $u\in\mathrm{Sp}(1)$.
@@
Since $C_{u}$ is linear for every $u\in\mathrm{Sp}(1)$, it coincides with its differential. We also note that the following identities hold
$\operatorname{Re}(u\alpha u^{-1}\overline{u\beta u^{-1}})$ $=\operatorname{Re}(\alpha\overline{\beta})$
$|1-(uqu^{-1})^{2}|$ $=|1-q^{2}|$
$|uqu^{-1}|$ $=|q|,$
for every $q\in\mathbb{B}$, $\alpha,\beta\in\mathbb{H}$ and $u\in\mathrm{Sp}(1)$. These remarks and the first formula for $G$ in Theorem 6.1 imply the claimed invariance. From this, the representation formula is just a restatement of the $C_{u^{-1}}$-invariance. ∎
@@
$\Omega_{q}(\alpha,\beta)=\frac{\operatorname{Im}\big{(}(1-q^{2})^{-1}(\alpha-q\alpha q)(\overline{\beta-q\beta q})(1-\overline{q}^{2})^{-1}\big{)}}{(1-|q|^{2})^{2}},$
for every $q\in\mathbb{B}$ and $\alpha,\beta\in\mathbb{H}$. In particular, $\Omega$ satisfies the following properties.
@@
The expression for $\Omega_{q}$ in the statement is a direct consequence of the definition of $\Omega$ and Theorem 5.7.
[…]
Finally, for a given $I\in\mathbb{S}$, the restriction of $\Omega$ to the tangent bundle $T\mathbb{D}_{I}$ is obtained by considering the values
@@
The equivariance formulas stated are easily obtained by substitution on the expressions for $H$ and $\Omega$ given in Theorems 5.7 and 6.7, respectively. ∎
@@
With the previous results at hand, we compare the $\mathrm{Sp}(1,1)$-invariant metric $\widehat{G}$ and a couple of well known geometric quaternionic structures with our slice geometric ones. For simplicity, we will discuss the $4$-dimensional case which is clearly enough for our purposes.
In the first place, we have the notion of a hyper-Kähler structure. On a $4$-dimensional manifold, such structure is given by a Riemannian metric and three integrable complex structures orthogonal for the metric and satisfying the usual quaternionic commutation relations. Further details on their definition and properties can be found in [2]. As stated in Theorem 14.13 of the latter reference, Berger proved that every hyper-Kähkler manifold is Ricci flat. However, the metric $\widehat{G}$ is known to have strictly negative sectional curvature (see [15]) and so strictly negative Ricci curvature. Hence, from Theorem 5.1 it follows that there is no hyper-Kähler structure on $\mathbb{B}$ that is invariant under the $\mathrm{Sp}(1,1)$-action by non-regular Möbius transformations.
In second place, there is the notion of quaternion-Kähler manifold. We refer again to [2] for further details. One possible definition, in the $4$-dimensional case, is obtained by requiring the given Riemannian manifold to have holonomy $\mathrm{Sp}(1)\times\mathrm{Sp}(1)$. This is in fact the case for the metric $\widehat{G}$, since the holonomy and the isotropy coincide for Riemannian symmetric spaces (see [15]). Hence, $\mathbb{B}$ with the metric $\widehat{G}$ is quaternion-Kähler. It is also known, as discussed in section D from Chapter 14 of [2] and in page 92 of [14], that every quaternion-Kähler manifold carries three locally defined almost-complex structures and three locally defined $2$-forms. The latter allow to construct a globally defined $4$-form. However, in the case of $\mathbb{B}$ with the metric $\widehat{G}$ the locally defined objects cannot be extended globally. In particular, the quaternion-Kähler structure obtained from the $\mathrm{Sp}(1,1)$-action does not yield associated globally defined $2$-forms.
In conclusion, the $\mathrm{Sp}(1,1)$-action by non-regular Möbius transformations fails to provide a notion of Kähler structure on $\mathbb{B}$ as some sort of $2$-form involving $\mathbb{H}$ that could also be nicely related to a corresponding Riemannian metric.
However, the slice regular Kähler structure $\Omega$ does provide a natural $2$-form with the Kähler-like properties obtained in Theorem 6.7. Furthermore, by construction, $\Omega$ exhibits as much invariance under $\mathcal{M}(\mathbb{B})$ as the latter allows considering that it is not a group. Since, $\Omega$ is $\operatorname{Im}(\mathbb{H})$-valued we can always consider the components with respect to an orthonormal basis of $\operatorname{Im}(\mathbb{H})$ to obtain three real-valued $2$-forms. The slice Hermitian metric $H$ and the slice Riemannian metric $G$ have similar properties as stated in Theorems 5.7 and 6.1, respectively, and their corollaries. More precisely, they both exhibit as much invariance as $\mathcal{M}(\mathbb{B})$ allows.
Also, a certain degree of invariance for all three slice geometric structures has been obtained in Corollaries 6.6 and 6.10. The latter have provided the sort of Representation Formulas found in the theory of slice regular functions, but now for geometric structures on $\mathbb{B}$ built from regular Möbius transformations.
Finally, the slice Riemannian metric $G$ and the slice Kähler form $\Omega$ can be considered nicely related in a similar way as it occurs in the classical complex case. The reason is that $G$ and $\Omega$ are the (quaternionic) real and imaginary parts, respectively, of the slice Hermitian metric $H$. With all three of them built out of the same principle: regular Möbius transformations.
@@ Acknowledgment
This research was partially supported by SNI-Conahcyt and by Conahcyt Grants 280732 and 61517.
Data Availability
Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Competing interests
The author has no competing interests to declare that are relevant to the content of this article.